a rectangle is 8m longer than it is wide. Its perimeter is 56 . Find its lenght and width
Answer by addingup(3677)

L = W+8
and
2L+2W = 56 substitute for L, per above:
2(W+8)+2W = 56
2W+16+2W = 56
4W = 40
W = 10
------------------------
Check:
L = 10+8 = 18
2(18)+2(10) = 36+20 = 56 Correct
:
John